# Argumentation Computation with Large Language Models : A Benchmark Study

### v1

@@ 5.2 Main Results
5.3 Errors in Computing Grounded Labelling
In this section, we summarize the errors and failures encountered by the LLMs when computing the grounded labelling.
@@ 5.4.3 Indefensibility
5.5 Effect of Explanations
In this section, we examine the effect of explanations on performance and the generalization capability of the model when applied to AAF with more arguments. To demonstrate the efficacy of explanations in data generation, as a comparison model, we additionally trained models, namely Llama3-noexp and Qwen2-noexp, on datasets without explanations, where the output text contained only the answer. To assess generalization capability across varying numbers of arguments, we generate additional test datasets spanning from 26 to 35 in number, each comprising 100 samples.
@@ 5.5.2 Generalization Stage
In the generalization stage, significant differences emerge between models trained with and without explanations. Models with explanations maintain relatively high accuracy, though performance gradually declines as complexity increases. Conversely, models without explanations experience a sharp drop in accuracy, sometimes falling to 20%, indicating poor generalization. This disparity underscores the importance of explanations in facilitating the acquisition of underlying logical structures, enhancing model robustness and adaptability to novel, complex scenarios. These findings emphasize the need for incorporating explanatory components in training to improve model resilience and generalization.
6 Related Work
Within the domain of formal logical deduction using language models, the primary focus has predominantly been on first-order logic Tian et al. (2021). Works such as RuleTaker Clark et al. (2020) and FOLIO Han et al. (2022) exemplify the application of artificial intelligence in deducing logical conclusions from structured premises, thereby demonstrating models’ capacity for formal logical reasoning. However, argumentative reasoning, which involves both logical reasoning and graph comprehension, has been largely overlooked Guo et al. (2023).
Previous research has attempted to integrate argumentation frameworks or with neural networks for interpretation Potyka (2021) and semantics computation Craandijk and Bex (2022, 2020). For instance, the work on interpreting neural networks as quantitative argumentation frameworks Potyka (2021) explores how neural networks can be understood through the lens of argumentation. Similarly, the AGNN Craandijk and Bex (2020) investigates the application of deep learning techniques to abstract argumentation semantics. Additionally, research in Craandijk and Bex (2022) proposes enforcement heuristics for argumentation with deep reinforcement learning, aiming to enhance the decision-making process in argumentation frameworks. In Mileva et al. (2023), researchers aim to consolidate diverse approaches to learning argumentation semantics within a unified framework.
Recent research has integrated argumentation techniques with large language models to enhance reasoning capabilities in various tasks.  Gorur et al. (2024) demonstrates that LLMs like Llama-2 and Mistral outperform traditional models in Relation-Based Argument Mining by effectively identifying support and attack relations between arguments.  Chen et al. (2024) conducts a comprehensive investigation into the potential of LLMs, such as ChatGPT, for computational argumentation. Their study focuses on argument mining and generation tasks, evaluating these models in both zero-shot and few-shot learning scenarios across diverse datasets.  Freedman et al. (2024) proposes an innovative approach to augment the reasoning abilities of LLMs through the incorporation of argumentative reasoning, which leverages LLMs to construct argumentation frameworks. The integration of these methodologies holds significant promise for advancing the development of robust artificial intelligence systems capable of sophisticated argumentation and reasoning.
@@ 7 Conclusion
In this paper, we examine the computation of abstract argumentation frameworks utilizing large language models. Our approach integrates argumentation computation with the capabilities of LLMs, providing enhanced explainability over traditional graph neural network methods. Experimental results demonstrate that LLMs can effectively learn and execute extension computation tasks. Key findings highlight the significant role of explanations in enhancing model performance and generalization. The self-explanatory capacities of LLMs address transparency challenges typically associated with neural networks. This research advances the understanding of LLMs in argumentation computation and paves the way for further investigation into defeasible reasoning and complex decision-making processes.
@@ Appendix A Chain-of-Thought Baseline
Prompting generally results in models making predictions with lower accuracy compared to models that have been fine-tuned with substantial training data Zhang et al. [2023]. To validate this observation and establish a baseline method, we design a Chain-of-Thought (CoT) pipeline Wei et al. [2022] that computes extensions based on a labelling algorithm. The argumentation computation is divided into multiple steps, framed as a reasoning task.

### v2

@@ 6.9.3 Indefensibility
6.10 Error Analysis of the GNN Baseline
Figure 9: A failure case of AGNN: The model incorrectly predicts the grounded extension as the empty set, while the correct answer is {4, 7}.
To compare the error patterns of GNNs and our LLM-based approach, we analyze the failure cases in AGNN [ijcai2020p231]. Our in-depth examination reveals a prevalent failure mode that does not exist in LLMs: in the vast majority of incorrect cases, the model fails to infer a valid grounded extension and instead incorrectly predicts the empty set, even when a correct extension could be easily found. An example of such a failure is shown in Figure 9. Because GNNs learn from data, the reasoning process behind their predictions lacks transparency, making it difficult to interpret intuitively and thereby reducing our confidence in the model’s predictions in these specific failure cases.
7 Empirical Study on Process Explanations
In this section, we examine the effect of explanations on model performance and generalization capability. We find that within the existing data distribution ($n\leq 25$), explanations have only a marginal impact on performance. While when applied to AAFs with more arguments, which we referred to as the generalization stage, models trained with explanations demonstrate stronger generalization ability. Additionally, we observe that human-annotated explanations further enhance the diversity of model-generated explanations, leading to better explanatory quality, which is important for human-AI interaction in real-world applications.
7.1 Role of Explanations in Generalization
To demonstrate the efficacy of explanations in data generation, as a comparison model, we additionally trained models, namely Llama3-noexp and Qwen2.5-noexp, on the datasets without explanations, where the output text contained only the answer. To assess generalization capability across varying numbers of arguments, we generate additional test datasets spanning from 26 to 35 in number, each comprising 100 samples.
@@ 7.1.2 Generalization Stage
In the generalization stage, significant differences emerge between models trained with and without explanations. Models with explanations maintain relatively high accuracy, though performance gradually declines as complexity increases. Conversely, models without explanations experience a sharp drop in accuracy, sometimes falling to 30%, indicating poor generalization. This disparity underscores the importance of explanations in facilitating the acquisition of underlying logical structures, enhancing model robustness and adaptability to novel, complex scenarios. These findings emphasize the need for incorporating explanatory components in training to improve model resilience and generalization.
7.2 Ablation Study
Model Exp. Component $_{\textsf{grd}} ACC$$_{\textsf{com}}$
Qwen2.5-SFT None 0.68 0.56
Llama3-SFT None 0.58 0.42
Qwen2.5-SFT Q 0.72 0.59
Llama3-SFT Q 0.66 0.53
[…]
Table 5: Ablation study on explanation components.
[…]
7.3 Effect of Human Feedback
Figure 11: Participant preferences for process explanations.
In the subsequent evaluation phase, we invited non-expert participants to perform pairwise preference tests, comparing explanations generated by four different models: Qwen2.5-7B, GPT-4o, Qwen2.5-SFT, and Qwen2.5-RLHF. For each of 50 randomly selected questions from the test set, 10 participants were asked to choose the explanation they found most understandable or helpful, providing a human-centered evaluation of explainability. Note that model outputs may contain errors, and participants are instructed to take these into account as part of their evaluation. The results are summarized in Figure 11. Our findings suggest that integrating human feedback via RLHF can substantially improve the explainability of model-generated decisions, making them align more closely with human reasoning and teaching practices.
8 Limitations and Future Work
Our training approach primarily focuses on abstract argumentation, enabling LLMs to resolve conflicts in a more logical manner. However, we do not explore the process of grounding real-world scenarios to abstract models. Mapping complex real-world situations to arguments and attack/support relations is another important issue, where fairness and representational bias may arise [chen2024credible, cramer2019empirical]. Encoding inductive biases into formal logical structures can directly impact the applicability of abstract argumentation in LLMs. Additionally, argumentation graphs derived from complex real-world situations may exhibit greater structural complexity, and our current data distribution may not comprehensively cover such edge cases.
There are two promising directions to address these key challenges. The first is to train models that can map conflict-rich real-world scenarios to formal representations, thereby improving the connection between practical applications and formal methods. The second is to leverage richer argumentation models to enhance the model’s ability to handle uncertain reasoning, including probabilistic settings [hunter2021probabilistic] and quantitative argumentation frameworks [potyka2021interpreting].
@@ 9 Conclusion
In this paper, we examine the computation of abstract argumentation frameworks utilizing large language models. Our approach integrates argumentation computation with the capabilities of LLMs, providing enhanced explainability over traditional graph neural network methods. Experimental results demonstrate that LLMs can effectively learn and execute extension computation tasks. Key findings highlight the significant role of explanations in enhancing model performance and generalization. The self-explanatory capacities of LLMs address transparency challenges typically associated with neural networks. This research advances the understanding of LLMs in argumentation computation and paves the way for further investigation into defeasible reasoning and complex decision-making processes. Our future works include exploring how LLMs can adapt to dynamic and incomplete argumentation frameworks, where arguments and relationships evolve or are missing, would enhance their applicability to real-world scenarios like legal reasoning and ethical decision-making.
Acknowledgments
This research was partially supported by the National Natural Science Foundation of China (62576309).
@@ Appendix A Chain-of-Thought Baseline
Prompting generally results in models making predictions with lower accuracy compared to models that have been fine-tuned with substantial training data [zhang2023large]. To validate this observation and establish a baseline method, we design a Chain-of-Thought (CoT) pipeline [NEURIPS2022_9d560961] that computes extensions based on a labelling algorithm. The argumentation computation is divided into multiple steps, framed as a reasoning task.
@@ A.0.2 Complete Semantics
Figure 14: Examples of two human annotations of explanation.
Appendix B Integrating Human Feedback
For human-annotated explanations, due to the limited availability of expert annotators, we focus on explanations for the computation of the grounded extension and annotate a subset of high-quality data. For each number of arguments, we annotate 50 distinct problems, resulting in a total of 1,000 training examples. For RLHF training, we further fine-tune the SFT-trained model using the GRPO algorithm [shao2024deepseekmath] within the TRL [vonwerra2022trl] framework, leveraging the collected human-annotated explanations as feedback. Figure 14 presents two examples of human annotations. To ensure the model outputs only the extension (rather than the label as in SFT), we manually add a prompt at the end requesting a JSON object containing the extension set. The reward function is defined as follows: a score of 1 is assigned if the extension is completely correct, and 0 otherwise.